OpenStudy (anonymous):

solve using elimination 2x+6y=5 x+2y=2

OpenStudy (anonymous):

2x + 6y = 5 2x + 4y = 4 (multiply second equation by 2) subtraction second equation from first 2y = 1 y = 1/2 plug back into equation 2 (original) x + 2y = 2 x + (2)(1/2) = 2 x + 1 = 2 x = 1 x = 1 y = 1/2